OpenStudy (anonymous):

In a hydrogen atom, the electron in orbit around the proton feels an attractive force of about 8.82x10^–8N. If the radius of the orbit is 4.24x10^–11m, and has a mass of 9.10939x10^–31, what is the frequency?Answer in revolutions/second.

OpenStudy (anonymous):

Knowns: m=9.10939x10^–31 kg attracted force (Fc)=8.82x10^–8N radius=4.24x10^–11meters

OpenStudy (anonymous):

Using \(F_c =m \frac{v^2}{r}\), and \(f=\frac{v}{2 \pi r }\) sub all the known in and you should get v.